OpenStudy (anonymous):

Consider the equation y= -2x+5 Create a table of five ordered pairs that satify the equation. What is the y-intercept of the equation? What is the x-intercept of the equation?

OpenStudy (anonymous):

did they give you specific values of x to use, or can you use any?

OpenStudy (anonymous):

a. y-intercept (5, 0); x-intercept -5/2,0 b. y-intercept (0, 5); x-intercept 0,5/2 c. y-intercept (5, 0); x-intercept 5/2,0 d. y-intercept (0, 5); x-intercept 5/2,0

OpenStudy (anonymous):

That's really about it

OpenStudy (anonymous):

ok. To find the y-intercept, plug in 0 for x and solve for y. y = -2x + 5 y = -2(0) + 5

OpenStudy (anonymous):

to find the x, do the opposite. plug in 0 for y and solve for x y = -2x + 5 0 = -2x + 5

OpenStudy (anonymous):

y-intercepts are always (0, number) x-intercepts are always (number, 0)

OpenStudy (denisaboichuk):

b